OpenStudy (anonymous):

Medal and fan! Given that the area of B=5, integral of (-2,1) f(x)dx= 2, and integral of (1,8) f(x) dx = -4. 1) integral of (-2,6) f(x) dx 2) Integral of (1,6) f(x) dx 3) integral of (6,8) f(x) dx 4) integral of (-2,8) [absolute value of f(x)] dx 5) integral of (-2,8) f(x) dx

OpenStudy (loser66):

I don't understand the problem. Is it \(\Large \int_{-2}^{1} f(x)dx =2\\ \Large \int_{1}^{8} f(x)dx =-4\) and then???

OpenStudy (anonymous):

yes, then solve the 5 exercises with the same information above. I think it has to do with integrating even/odd functions? because I am correcting

OpenStudy (loser66):

oh, I got it

OpenStudy (loser66):

that is the given information and you have to answer 5 questions bellow, right? with the whole Area is 5, right?

OpenStudy (anonymous):

Yes, I think so

OpenStudy (loser66):

It's not hard. I do the first one for you. You can imitate for the leftover, deal?

OpenStudy (anonymous):

ok

OpenStudy (loser66):

first off, the whole area is 5 it means after finding the total area is 5-->\[ \int_{-2}^{1} f(x) dx +\int_{1}^{6} f(x)dx=\int_{-2}^{6}f(x)dx =5\]

OpenStudy (loser66):

but it doesn't help much, I do the next one for you as sample From part 1) you have \[\int_{-2}^{1} f(x) dx +\int_{1}^{6} f(x)dx=\int_{-2}^{6}f(x)dx =5\\\] \[\rightarrow\int_{1}^{6} f(x)dx=\int_{-2}^{6}f(x)dx -\int_{-2}^{1} f(x) dx = 5-2 =3 \]

OpenStudy (loser66):

can you do the rest?

OpenStudy (anonymous):

I did #5 and only got to \[\int\limits_{-2}^{8} f(x) dx = \int\limits_{2}^{1} f(x) dx + \int\limits_{1}^{8} f(x) dx\] The only thing is that I don't know how you get the 5 from the first and the 3 from the second exercise?

OpenStudy (loser66):

the given information is " Given the area of B is 5"

OpenStudy (anonymous):

I get where you got the 5-2 from the info, but then for the 5th exercise I did it would be ? - (-4)?

OpenStudy (loser66):

wait, I' ll scan the explanation for you

OpenStudy (anonymous):

Ok yes, the other one I'm having total difficulty is the one with the absolute value exercise, a little hint for that one? :)

OpenStudy (loser66):

just take positive value, it means 4 instead of -4

OpenStudy (anonymous):

and for \[\int\limits_{6}^{8} f(x)dx = \int\limits_{?}^{?} f(x) dx - \int\limits_{1}^{8} f(x)dx = ? -( -4) ?\]

OpenStudy (loser66):

you have 1to 6 and 1to 8 subtract them to get 6 to8

OpenStudy (loser66):



OpenStudy (anonymous):

Alright thanks a lot! Great explanations.